Operations Report – January 30th


Crew 274 Operations Report 1/30/2023

SOL: 1

Name of person filing report: Alexis Lojek

Non-nominal systems: Oven light

Notes on non-nominal systems: Oven light still needs a new bulb. We previously reported that the specifications of the oven light bulb type were emailed to mission support by Crew 272 Engineer.


Spirit rover used: Yes

Hours: 207.5

Beginning charge: 100%

Ending charge: 91%

Currently charging: No

Opportunity rover used: Yes

Hours: 112.2

Beginning charge: 100%

Ending charge: 91%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 219.5

Beginning charge: 100%

Ending charge: 91%

Currently charging: No

Perseverance rover used: Yes

Hours: 254.2

Beginning charge: 100%

Ending charge: 87%

Currently charging: Yes

General notes on rovers: All four rovers were used on EVA during second rover training mission to Marble Ritual. The chargers were set aside upon return to the Hab.

Summary of Hab operations: It was noted that the oven light was not working and that Crew 274 had also reported the same. The oven currently has no bulb in it, but the bulb cover is still present. No showers were taken today.

WATER USE: 46 gallons

Water (static tank): 445 gallons remaining

Static tank pipe heater (On or off): On

Static tank heater (On or off): On

Toilet tank emptied: No

Summary of internet: Nominal

Summary of suits and radios: Suits 1 & 4 were not utilized during EVA suit training due to Suit 1 blowers not turning on and Suit 4 having a crossthreaded coupling. The crossthreaded coupling for the vent fans on Suit 4 was fixed after the EVAs were completed. Suit 1 remains an issue and has not yet been troubleshot. Suits 2, 3, 5, & 6 were utilized during both the first and second EVAs. Helmets were cleaned before and between EVAs and different helmets were utilized for different crew members.

Summary of GreenHab operations: GHO Hines watered plants this morning at 0830-0930 am.

WATER USE: 12 gallons used.

Heater: On

Supplemental light: 4 hours

Harvest: None

Summary of ScienceDome operations: GreenHab Officer Tyler Hines continued setting up his equipment in the ScienceDome.

Dual split: Unknown, request MDRS Support guidance on what this is.

Summary of RAM operations: No operations were conducted in the RAM.

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support: What is the “dual split?” in the ScienceDome? Should we attempt to troubleshoot EVA 1?

Copyright © The Mars Society. All rights reserved. | Main Site